OpenStudy (anonymous):

please help, stuck on this problem. find difference of f, f(x+h)-f(x) over h f(x)=5x^2-9x+8

OpenStudy (anonymous):

f '(x)=10x-9

OpenStudy (anonymous):

I think the purpose is to evaluate the limit of the difference quotient instead of simply differentiating.

OpenStudy (anonymous):

\[\frac{5x^{2}+10xh+5h^{2}-9x-9h+8-5x^{2}+9x-8}{h}=\frac{10xh+5h^{2}-9h}{h}=10x-9\]

OpenStudy (anonymous):

when you take the limit in the last step of course.

OpenStudy (anonymous):

To clarify what he did: \[\lim_{h \rightarrow 0} \frac{f(x + h) - f(x)}{h}\]

OpenStudy (anonymous):

should it be 10x-9h? or just 10x-9?

OpenStudy (anonymous):

the limit is the derivative, 10x-9

OpenStudy (anonymous):

thank you both for your help! much appreciated